HUBBLE SPACE TELESCOPE DAILY REPORT #2121 PERIOD COVERED: 0000Z (UTC) 05/11/98 - 0000Z (UTC) 05/12/98 Daily Status Report as of 132/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was to be used to probe the dark matter distribution of NGC-7212 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. This proposal was not executed due to NICMOS being in the suspend state. 1.2 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/MA2 7647 (STIS NUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spectrograph (MA2) was used to make near-ultraviolet (NUV) flat calibration images. This program will obtain NUV-MAMA flat-field observations with the D2 lamp for the construction of on-orbit D-flats for G230M. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/1/2/3 7902 (Photometric Monitoring, Photometric Monitoring-II) The Near Infrared Camera (Cameras One, Two and Three) was to be used to make observations of absolute standard, solar analog star P330E-Cal as part of a photometric monitoring program. Observations were made at several different wavelengths spanning the NICMOS wavelength range in each camera NIC/1, NIC/2, and NIC/3 to monitor the photometric stability of the NICMOS cameras as a function of wavelength. This proposal was not executed due to NICMOS being in the suspend state. 1.5 Completed Two Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.6 Completed STIS/CCD 7645 (STIS FUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spaectrograph (CCD) was used to make far-ultraviolet flat calibration images. This program will obtain FUV-MAMA flat-field observations with the Kr lamp for the construction of on-orbit D-flats for G140M. A dark calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.7 Completed Four Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) The Near Infrared Camera (Camera One) was to be used to make observations of the Seyfert Galaxy NGC-4388. These data will allow us to determine if the Seyfert Type 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. This proposal was not executed due to NICMOS being in the suspend state. 1.9 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.10 Completed STIS/CCD 7438 (Spatially-Resolved Spectroscopy of the UVX in Elliptical Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make spatially resolved far-ultraviolet observations of the large elliptical galaxy NGC-3115. The prominent far-UV upturn {UVX} in the spectra of elliptical galaxies originates from stars in their old populations which have evolved beyond the He-flash at the tip of the red giant branch. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was to be used to make pure parallel spectrographic observations of sky regions when another instrument is prime. This proposal was not executed due to NICMOS being in the suspend state. 1.12 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed NIC/2 7877 (Infrared Luminous Stars & Stellar Population in Nearby Giant Elliptical Galaxies) The Near Infrared Camera (Camera Two) was to be used to make observations of infrared luminous stars in the nearby giant elliptical galaxy NGC-5128. These data will yield the first infrared luminosity functions for the stellar populations of giant ellipticals as well as unique optical/infrared color-magnitude diagrams for individually resolved stars. This proposal was not executed due to NICMOS being in the suspend state. 1.14 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.15 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was to be used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. This proposal was not executed due to NICMOS being in the suspend state. 1.16 Completed NIC/1/2/3 7857 (Investigating the Missing Link Between Disks Around Pre-Main Sequence and Main Sequence Stars) The Near Infrared Camera (Cameras One, Two and Three) was to be used to make observations of stars HD100546 and HD106797. Various observations and models of the circumstellar disk around the Main Sequence IR excess star Beta Pictors tend to show that the optically thin dust disk is the result of destruction processes among planetesimals. This proposal was not executed due to NICMOS being in the suspend state. 1.17 Completed STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed NIC/1/2/3 7844 (H-2 Observations of the Galactic Center Circumnuclear Ring) The Near Infrared Camera (Cameras One, Two and Three) was to be used to make observations of the galactic center circumnuclear ring. These data will be used the distribution of molecular hydrogen in the circumnuclear ring that engulfs the inner two parsecs of the galaxy. This proposal was not executed due to NICMOS being in the suspend state. 1.19 Completed STIS/MA1/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.20 Completed WF/PC-2 7470 (A Snapshot Survey of Galactic Globular Clusters) The WF/PC-2 was used to make snap shot survey observations of the crowded center of the galactic globular cluster NGC-6342. The resulting color-magnitude diagram will include all stars down to about one magnitude below the main-sequence turnoff. The observations were completed as scheduled, and no problems were reported. 1.21 Completed NIC/2 7862 (A Search for Near Infrared H-2 Emission in Active Elliptical Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of active elliptical radio galaxy 3C264 to search for infrared H-2 region emissions. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: Using ROP SR-1, the SSR correctable EDAC counts were cleared four times. Per operations request (OR) 13751, NICMOS was recovered from suspend to operate at 131/205517Z. By the same OR, the PAM was moved to its home position, the buffer box temperature low limits were patched, and memory monitors were restored by 131/2146Z. The proposals described in 1.1, 1.4, 1.8, 1.11, 1.13, 1.16, and 1.18 were affected by the suspend. NICMOS was recovered just before the proposal described in 1.21. That proposal operated nominally. Per ROP RD-7, ESTR-2 was re-conditioned at 132/0512Z and ESTR-3 was re-conditioned at 132/0941Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH